Math 291 Presentation Lab Due the day that you present (via email)

Instructions: Turn in your lab by emailing it to fagerstrom@mnstate.edu. You should email both your raw TEX (.tex) file and your compiled Beamer document (in .pdf form).

You will be graded on both your raw TEX code and the accuracy of your compiled document. Name your file in the format of JaneDoePresentation.tex

• Your presentation is worth 40 points overall.

• 20 points is for the presentation in class itself, which is detailed on the separate rubric.

• 20 points is for the Beamer file and its original .tex code, which is detailed below.

Your Beamer file for the presentation must include the following elements. Note that everything that is in your Beamer presentation should be referred to during your presentation at some point, so make sure that these elements are connected in some way.

• You must use Beamer.

• There must be math equations.

• There must be math graphs (graphs of functions, 3D graphs, trees or other graphs using nodes, etc.)

• Note that both of the above are plural. You should also have more than a total of four of them, although how they are split will be determined by your topic.

• There must be at least one picture (.jpg or similar). What is it is up to you - a picture of you on the title page, the MSUM logo, a picture of the physics or biology lab that relates to your work, etc. — whatever is appropriate for your topic.

• There must be a title page and a table of contents.

• There must be some sort of introduction to your topic, at least two sections in the main body, and some sort of conclusion or other wrap-up page.

• The length should be appropriate to accompany a 5-10 minutes presentation.

Additional comments

• I would actually prefer presentations in the 5-7 minute range, so that we can easily fit 6 per day for two class days (and then not have to use the final exam time).

• If you want to use this presentation as a dry run for a presentation in another class or project, or a re-presentation, that is fine, just make sure that you let both me and your other professor know that.

• The content is very open-ended: Make it connect to math in some way. It can be a lesson plan for a high school level class, a physics or other science topic that connects to math enough that the requirements above make sense, a presentation of problems from your most recent Calculus II lab, even math jokes. It’s up to you.